A medallion made of nickel (coefficient of linear expansion = 1.5 x 10-5/K) has a diameter of 0.1m at 20oC. What will be its diameter at 35oC?


1
Expert's answer
2022-01-18T10:24:02-0500

Linear expansivity"(\alpha)={l_2 - l_1 \over l_1(\theta_2 -\theta_1)}"

"\alpha=1.5×10^{-5}"

"l_1=0.1m"

"l_2=?"

"\theta_1=20°C=293K"

"\theta_2=35°C=308K"


"l_2= \alpha l_1(\theta_2 - \theta_1)+l_1"

"l_2=1.5\u00d710^{-5}\u00d70.1(308-293)+0.1"


"l_2=0.1000225m"